OpenStudy (anonymous):

Solve by completing the square. Can't use any other method. 3x^2-6x=2

OpenStudy (anonymous):

3x^2-6x=2 3(x^2-2x)=2 3(x^2-2x+1)=2+3 3(x-1)^2=5 (x-1)^2=5/3 (x-1)=±√(5/3) x =1 ± √(5/3)